# Uncertainty-Aware Critic Augmentation for Hierarchical Multi-Agent EV Charging Control

### v4

@@ A Reproducibility
A.1.4 Baseline Settings.
To evaluate the performance of the proposed HUCA model, we compared it against several established baseline approaches:
  • DDPG (Deep Deterministic Policy Gradient) [13]: An actor-critic algorithm designed for continuous control in single-agent environments.
  • IQL (Independent Q-Learning) [22]: A decentralized multi-agent reinforcement learning approach where each agent independently learns its own Q-function.
  • VDN (Value Decomposition Networks)[21]: A centralized value-based reinforcement learning method used in multi-agent systems.
  • MADDPG (Multi-Agent Deep Deterministic Policy Gradient) [14]: Extends DDPG to multi-agent settings by incorporating centralized training and decentralized execution.
A.2 State representations of high-level and low-level MDPs.
A.2.1 High-level Agent.
A state $s_{h}$ in the MDP optimizied by the high-level agent includes three main information $s_{h}=(I^{e}_{t},I^{v}_{t},I^{l}_{t-1})$ at current time slot $t$. The details are as follows:
  • $I^{l}_{t-1}$ includes the average critic value (from the action-value function) and the standard deviation of all low-level agents from the previous time slot.
A.2.2 Low-level Agent.
For a low-level agent $c_{i}\in\mathcal{C}$, the state $s^{i}_{l}$ observed at time slot $t$ can be described as comprising two main information $s^{i}_{l}=(I^{v_{j}}_{t},I^{h}_{t})$. Details are:
A.3 Optimization of HUCA.
We optimize the HUCA framwork using minibatches of samples from the replay buffers. For the high-level agent, given a minibatch of $\mathcal{M}$ samples from the replay buffer $\mathcal{D}_{h}$, the training loss of the critic network is rewritten as:
(A.2) $\mathcal{L}(\theta^{Q}_{h})=\frac{1}{\mathcal{M}}\sum_{m}\Bigl{(}Q_{h}(s^{m}_{h},a^{m}_{h})-y^{m}_{h}\Bigr{)}^{2}.$
The actor network is updated using the sampled policy gradient as:
(A.3) $\nabla_{\theta^{\mu}_{h}}J(\mu_{h})\approx\frac{1}{\mathcal{M}}\sum_{m}\nabla_{\theta^{\mu}_{h}}\mu_{h}(a^{m}_{h}|s^{m}_{h})\nabla_{a_{h}}Q_{h}(s^{m}_{h},a^{m}_{h}).$
Similarly, the critic and actor networks of each low-level agent $c_{i}$ are also optimized using minibatches.
B Limitation and Future Work
Despite the strong performance of HUCA and its consideration of multiple dynamic factors, we have not yet accounted for user willingness to adjust their charging preferences to supply energy to the building. A promising direction for future work is to incorporate user willingness, along with the buy-and-sell energy behaviors between users and the system, into our operational framework. User willingness could be integrated into the low-level actor-critic networks of HUCA. In our current design, the departure uncertainty term is used to set a lower confidence bound for the action-value function of each low-level agent. When incorporating user willingness, an incentive mechanism could be developed to estimate the upper confidence bound of the action-value function, as user willingness and incentives might increase the effectiveness of charging control.
This would not only help identify which EV users have greater flexibility in participating in energy control but also allow the system to evaluate both the lower and upper bounds of the potential quality of low-level actions. This approach could improve the optimal charging and discharging power control, enhancing both user satisfaction and the overall effectiveness of the system.

### v5

@@ 5.1 Dataset and Experimental Setup
5.2 Comparison Results.
As shown in Table 3, except for the ”OPT” method (best-case result), HUCA achieves the lowest penalty and total cost without relying on future information (such as future energy prices, EV charging requests, and actual EV departure times) for charging control across both departure scenarios. In addition, its total cost is comparable to OPT, manifesting its effectiveness in minimizing costs. While the SoC fulfillment, maintenance, and user satisfaction scores of HUCA are not the best, its performance exceeds the median of the baselines in most cases (highlighted in gray). In contrast, although DDPG and IQL achieve higher SoC fulfillment under different numbers of charging piles, they incur substantially higher total costs than HUCA (by 20-80%).
Combining the above observations, HUCA actually strikes the best balance between the trade-off of electricity costs and fulfilling user demands by introducing the hierarchical control and the uncertainty-aware critic augmentation.
Among multi-agent models, VDN and MADDPG come closest to HUCA in terms of cost. However, user satisfaction scores of MADDPG are much lower than those of HUCA. While VDN achieves higher SoC fulfillment than HUCA with 20 charging piles, its maintenance scores are significantly lower. In some cases, the maintenance scores of MADDPG and VDN fall into negative values, indicating that EVs primarily serve as power providers and return with lower SoCs than at arrival, leading to additional expenses for users.
[…]
Uncertain Departure Scenario
CP Method Penalty ($\downarrow$) Ful. ($\uparrow$) Main. ($\uparrow$)
10 HUCA 355.84 40.77 45.40
w/o C.A. 2015.62(▽466.4%) 62.49(▲53.3%) 124.01(▲173.1%)
w/o H. 2707.71(▽660.9%) 63.10(▲54.8%) 31.91(▽29.7%)
w/o Either 723.46(▽103.3%) 44.53(▲9.2%) -12.53(▽127.6%)
CP Method Penalty ($\downarrow$) Ful. ($\uparrow$) Main. ($\uparrow$)
20 HUCA 354.90 42.44 124.41
w/o C.A. 9362.87(▽2538.2%) 69.73(▲64.3%) 209.70(▲68.6%)
w/o H. 3461.37(▽875.2%) 58.61(▲38.1%) 7.69(▽93.8%)
w/o Either 618.46(▽74.3%) 53.26(▲25.5%) 43.02(▽65.4%)
5.3 Ablation Study and Uncertainty Impact.
We conduct an ablation study by removing the uncertainty-aware critic augmentation (w/o C.A.), the high-level agent (w/o H.), and both of them (w/o Either). The results are shown in Table 4 and Figure 4. In Table 4, removing either the critic augmentation or the high-level agent increases SoC fulfillment and maintenance scores (improvements of about 9–173%). However, this also results in a much larger rise in penalty costs (worsening by about 74–2538%). The increase in penalty costs far outweighs the gains in SoC fulfillment and maintenance when removing each component. On the other hand, removing both components results in a slight increase in costs but causes a significant drop in SoC maintenance scores. Figure 4 also shows that with more charging piles (CP=20), removing critic augmentation increases user satisfaction but also raises the total cost substantially (worsening by 154%) compared to the full HUCA model.
Figure 4: Ablation study on total cost and user satisfaction (CP=20).
These findings indicate that the hierarchical structure and critic augmentation in HUCA complement each other, and their combined use is essential for achieving the lowest costs while balancing energy supply between the building and EVs.
5.4 Case Study.
To examine how HUCA balances energy supply between the building and EVs, Figure 6 visualizes the control decisions over a single day for all EVs, which can be discussed in terms of charging and discharging behaviors:
Charging Behaviors. Two significant EV charging behaviors appear at 7:00 and 16:00 (red rectangle) but with different reasons. At 7:00, both electricity demand and prices for the building are minimal (Figure 6(a)), therefore HUCA focuses on charging the EVs to reach their target SoC. At 16:00, even though there is a high demand for building energy and elevated prices, which typically would lead to transferring energy from the EVs to the building to reduce costs, HUCA opts to charge the EVs because of their high charging urgency (Figure 6(b)).
Discharging Behaviors. A similar situation occurs at 8:00 (green rectangle); however, since the charging urgency is relatively low at this time, HUCA discharges the EVs to serve the high-usage building while simultaneously avoiding high-priced electricity usage.
These results manifest that HUCA effectively and dynamically balances energy supply between the building and EVs based on real-time information.
6 Conclusion and Future Work
In this paper, we propose HUCA, a novel hierarchical multi-agent framework designed for real-time charging control. To address practical dynamics and limitations, HUCA integrates hierarchical control and uncertainty-aware critic augmentation to adapt to dynamic factors, optimize charging decisions, and account for departure uncertainties. Experiments show that HUCA outperforms existing methods in both certain and uncertain departure scenarios. A case study further illustrates that HUCA effectively balances energy supply between the building and EVs using real-time information.
One line of the future work is to account for user willingness when adjusting charging preferences to supply energy to the building, e.g., with both user willingness and buy-and-sell energy behaviors in HUCA. Incorporating them can help improve the estimation of the confidence bound for the action-value function, leading to a more efficient charging control.